When 50.2g iron(III) oxide reacts with carbon monoxide, 31.7g iron is produced. What is the percent yield of the reaction?
1
Expert's answer
2012-10-31T08:21:20-0400
η = (mpr/mteor)*100%;

Fe2O3 + 3CO = 2Fe + 3CO2

n(Fe2O3) = 50.2/160 = 0.31375 mol
n(Fe) = 0.6275 mol
mteor = 0.6275 * 56 = 35.14 g
η = 31.7 / 35.14 * 100% = 90.2 %